kekeman:

Shane performed the following trials in an experiment. Trial 1: Heat 30.0 grams of water at 0 °C to a final temperature of 40.0 °C. Trial 2: Heat 40.0 grams of water at 10.0 °C to a final temperature of 40.0 °C. Which statement is true about the experiments? The same amount of heat is absorbed in both the experiments because the product of mass, specific heat capacity, and change in temperature are equal for both. The same amount of heat is absorbed in both the experiments because the heat absorbed depends only on the final temperature. The heat absorbed in Trial 2 is about 3,674 J greater than the heat absorbed in Trial 1. The heat absorbed in Trial 2 is about 5,021 J greater than the heat absorbed in Trial 1.

Vocaloid:

q = mC*(Tf- Ti) where m is the mass, C is the specific heat of water (4.186 J/gC), Tf is final temperature, Ti is initial temperature, and q is heat absorbed. calculate q for both trials and compare the values.

kekeman:

I think it is " The heat absorbed in Trial 2 is about 3,674 J greater than the heat absorbed in Trial 1."

Vocaloid:

can you show me your calculations? (or at least how you set up the equation?) I got something a bit different.

kekeman:

Lol i fr just guessed this stuff confuses me

Vocaloid:

starting with the equation: q = mC*(Tf- Ti) m is mass, C is specific heat, Tf is final temperature, Ti is initial temperature. so for trial 1: mass is 30g, C is 4.186 J/gC, Tf is 40C, Ti is 0C so the equation becomes q = 30(4.186)(40-0) can you try setting up the equation for trial 2 using this same logic?

Vocaloid:

you're given all the information you need in the problem. you just need to plug in the numbers into the right places. for trial 2: q = mC*(Tf- Ti) Trial 2: Heat 40.0 grams of water at 10.0 °C to a final temperature of 40.0 °C. so m = 40g, C = 4.186, Tf = 40, Ti = 10 so the equation becomes q = 40(4.186)(40-10)

Vocaloid:

so comparing trial 1 vs trial 2: trial 1: q = 30(4.186)(40-0) trial 2: q = 40(4.186)(40-10) notice anything about these two equations?

kekeman:

The same amount of heat is absorbed in both the experiments because the heat absorbed depends only on the final temperature.

Vocaloid:

it's true that the two equations are both the same value, but that's not the reason why. the total heat depends on mass, specific heat, and temperature change (that's why all three values are included in the equation) in our case, notice how both equations are 30 * 4.186 * 40. that means the product of mass, specific heat capacity, and change in temperature are equal for both.

kekeman:

Omg geez i am so dumb

kekeman:

So it would be " The same amount of heat is absorbed in both the experiments because the product of mass, specific heat capacity, and change in temperature are equal for both."

Vocaloid:

yup good

kekeman:

Sorry i am kinda bad at this information